OpenStudy (anonymous):

if lim (6n^2)/(200-4n+kn^2)=1/2, then k= x->inf

jimthompson5910 (jim_thompson5910):

x is nowhere to be found in that expression did you mean n instead of x?

OpenStudy (anonymous):

yea sorry

jimthompson5910 (jim_thompson5910):

that's ok

OpenStudy (anonymous):

he can help more than i can sorry

jimthompson5910 (jim_thompson5910):

Notice how the degree of the numerator (2) is equal to the degree of the denominator (also 2). So that means that as n approaches infinity, it will equal the ratio of the leading coefficients. In other words, you divide 6 by k to get 6/k, then you set this equal to the given limit 1/2 and solve for k so... 6/k = 1/2 6*2 = k*1 12 = k k = 12

OpenStudy (anonymous):

its ok thank you

OpenStudy (anonymous):

so pretty much the infinities cancel out

jimthompson5910 (jim_thompson5910):

in a sense, yes

OpenStudy (anonymous):

thanks, you are awesome. one more problem and im done for my quarter project

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

help me with my last problem plsss

jimthompson5910 (jim_thompson5910):

sure, go ahead and post